OpenStudy (anonymous):

limit[x-(x^2-1)^1/2] as x approach infinity

OpenStudy (experimentx):

rationalize it \[ \frac{(x - \sqrt{x^2 - 1})(x + \sqrt{x^2 - 1})}{x + \sqrt{x^2 - 1}}\]